Convert into a mole.
20 g of water
Convert into a mole.
20 g of water
 

Grade:12th pass

5 Answers

Harshit Singh
2 years ago
Dear Student

Given mass of water = 20 g
Molar mass of water (H2O) = (2 × 1) + 16 = 18 g
Mole of water= 20/18 =1.12 mole

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Molar mass of Water = 18 g
Given mass of Water = 20 g
No. Of moles = Given mass/ Molar mass
= 20/18 = 1.12 moles
 
Hope it helps!
Thank you ✌️
irshad hussain
2 years ago
we know the molecular mass of hydrogen is 1 and oxygen is 16.

we know the formula of water is H2O

H2O
=2×1+16
= 2+16
= 18g
So, one mole of water weight is 18g.

therefore no of mole = give mass/molar mass

so, 20/18 = 1.11 moles
 
irshad hussain
2 years ago
we know the molecular mass of hydrogen is 1 and oxygen is 16.

we know the formula of water is H2O

 

Molar mass of Water =H2O=2+16=18 g/mol 
So, 1 mole of water weighs = 18 g 
which implies, 18g of water = 1 mole 
Therefore, 20g of water =20/18=1.11 mole 
H2O
=2×1+16
= 2+16
= 18g
So, one mole of water weight is 18g.

therefore no of mole = give mass/molar mass

so, 20/18 = 1.11 moles
irshad hussain
2 years ago

 

The molar mass of Water =H2O=2+16=18 g/mol 
So, 1 mole of water weighs = 18 g 
which implies, 18g of water = 1 mole 
Therefore, 20g of water =20/18=1.11 mole